1. In 10 hours A walks 5 miles more than B does in the same time; & in 8 hours B walks 8 miles more than A does in 4 hours. How many miles does each walk per hour?
Answer by ewatrrr(24785)

r A's mph, s B's mph
D = rt
10r - 10s = 5mi
8s -4r = 8mi
............
r - s = 1/2
-r + 2s = 2
s = 2.5 mph,B's mph and r = 3mph, A's mph r+-+2.5+=+.5
And...checking
30mi - 25mi = 5mi